1. how do you convert 3/7 to a percentage? (round to the nearest tenth if needed)?
2. How do you convert 0.89 to a percentage?
3. How do you convert 5% into a fraction and reduced to lowest term?
4) 100 is 85% of ?

Answer by checkley77(12844)

convert 3/7 to a percentage?
3/7=.4 ANS.
--------------------------------
convert 0.89 to a percentage?
.89=89% ANS.
-------------------------------
convert 5% into a fraction and reduced to lowest term?
5%=.05=5/100=1/20 ANS.
--------------------------------
100 is 85% of ?
100/.85=117.647 ANS.
--------------------------------